# Quaternion Multi-focus Color Image Fusion

### v1

@@ Abstract
Multi-focus color image fusion refers to integrating multiple partially focused color images to create a single all-in-focus color image. However, existing methods struggle with complex real-world scenarios due to limitations in handling color information and intricate textures. To address these challenges, this paper proposes a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain. This framework introduces 1) a quaternion sparse decomposition model to jointly learn fine-scale image details and structure information of color images in an iterative fashion for high-precision focus detection, 2) a quaternion base-detail fusion strategy to individually fuse base-scale and detail-scale results across multiple color images for preserving structure and detail information, and 3) a quaternion structural similarity refinement strategy to adaptively select optimal patches from initial fusion results and obtain the final fused result for preserving fine details and ensuring spatially consistent outputs. Extensive experiments demonstrate that the proposed framework outperforms state-of-the-art methods.
@@
Multi-focus color image fusion, quaternion color image fusion, quaternion sparse decomposition.
publicationid: pubid: $0$–$00$ 2025 IEEE
@@ I Introduction
A single color image often fails to capture all objects in focus due to optical limitations of cameras [1]. When imaging scenes with objects at varying distances, only regions within the focal plane are focused while areas outside this plane exhibit defocus blur. To circumvent this limitation, multi-focus color image fusion (MCIF) techniques integrate multiple partially focused color images of the same scene using focus detection and adaptive fusion strategies to obtain an all-in-focus color image [2]. This technology has numerous applications across diverse domains including digital photography, preprocessing operations [3], and target recognition [4, 5].
Existing MCIF methods can be broadly categorized into deep learning-based and conventional methods. Currently, deep learning-based methods have extensively promoted the development of MCIF to enhance the ability of learning multi-scale features to distinguish focused pixels [6, 7]. Since it is impractical to obtain perfect ground truth (i.e., all-in-focus color images) for real-world scenes [8], many deep learning-based works heavily rely on synthetic training data with a large-scale training process. Unlike deep learning-based MCIF methods, conventional methods only depend on human prior knowledge to distinguish focused objects without the training phase [9].
Existing conventional MCIF methods often treat color image fusion as a direct extension of multi-focus grayscale image fusion techniques by channel-wise processing [10] or grayscale conversion [11]. For channel-wise processing, these MCIF methods simply treat color channels as three independent grayscale images and directly apply image fusion techniques to each channel to obtain the fused color image. For instance, Yang et al. [10] extended sparse coding techniques to MCIF by separately measuring focus levels in each channel and fusing sparse coefficients channel-by-channel. This may introduce spectral-spatial inconsistencies and color distortions in the fused color images [12, 13].
@@
For grayscale conversion, existing conventional MCIF methods first convert the input color images into grayscale images and perform intensity-driven focus detection [14, 11] to derive focus maps from grayscale inputs to guide the fusion of original color channels [15]. For instance, [16] averaged color channels into a grayscale image and obtained focus maps by multi-scale morphological filtering. [17] transformed the color image into the YUV color space and detected focus levels employing the zero crossing and canny edge detectors. Although these methods avoid explicit channel separation, they fail to detect multi-channel sharpness variations among the input color images and may produce inaccurate focus maps under complex real-world scenarios. This leads to ghosting artifacts or blurred boundaries in the fused results [15, 13]. Therefore, it remains a challenge to perform high-quality color image fusion in real-world applications. The limitations of conventional MCIF methods lie mainly in two key challenges: (1) small-scale focus detection failure: existing techniques struggle to distinguish and retain small focused regions within images containing extensive defocused areas based on grayscale focus detection [4]; (2) boundary artifacts: focus maps derived from intensity-based or edge-based focus detection often produce inaccurate results with visual artifacts introduced particularly in uncertain boundary regions (between focused and defocused areas) and low-gradient regions due to defocus spread effects [15, 18].
Recently, a quaternion higher-order singular value decomposition (QHOSVD) model [19] was developed to perform quaternion-domain multi-focus color image fusion by fully considering highly-correlated color information. QHOSVD treats three color channels as a unified entity of three imaginary parts in a pure quaternion matrix using quaternion representation [20, 21] and extends this representation into a third-order quaternion tensor. This effectively preserves cross-channel relationships during color image processing [22, 23, 24]. Taking advantage of higher-order singular value decomposition in the quaternion domain, QHOSVD extracted focus levels corresponding to each source color image and employed a simple patch-wise fusion strategy to ensure spatial consistency and color fidelity [19]. However, QHOSVD still fails in focus information preservation and artifact suppression in complex real-world scenarios due to defocus spread effects [18].
@@
To address these issues, we propose a quaternion multi-focus color image fusion (QMCIF) framework to produce high-quality fused color images. Our main contributions are presented as follows:
@@
  • We propose a QMCIF framework to perform high-quality color image fusion completely in the quaternion domain. QMCIF is flexibly extended from fusing only two color images at a time into fusing multiple color images simultaneously under various challenging scenarios.
  • To effectively measure the focus level in both small focused areas and uncertain boundary regions, QMCIF introduces a quaternion sparse decomposition model to jointly learn fine-scale image details and structure information of color images in an iterative fashion.
@@
  • To preserve color structure and detail information, QMCIF introduces a patch-wise quaternion base-detail fusion strategy to fuse multiple partially focused color images in detail-scale and base-scale individually.
@@
  • To balance the trade-off between focus information preservation and artifact removal, QMCIF further introduces a quaternion structural similarity refinement strategy to flexibly form the final fused results.
@@
  • Extensive experiments on various datasets demonstrate that QMCIF outperforms the state-of-the-art methods.
@@
The rest of this paper is organized as follows: Section II presents the preliminaries of quaternion algebra. Section III introduces our framework and extends our framework to multi-color image fusion. Section IV presents the experiments and comparisons. Finally, Section V gives the conclusions.
@@ II Preliminaries
${(\cdot)}^{\mathrm{T}}$$\overline{(\cdot)}$,${(\cdot)}^{\mathrm{H}}$, ${(\cdot)}^{-1}$ transpose, conjugate , conjugate transpose and inverse representation
@@
    $Frobenius$ norm: ${\|\dot{\bm{\mathrm{A}}}\|}_{F}={(\sum_{s=1}^{M}\sum_{t=1}^{N}{{|\dot{a}_{s,t}|}^{2}})}^{\frac{1}{2}}$
@@
Given a color image $\bm{\mathrm{I}}$ in real three-dimensional space, its quaternion representation is defined as:
$\begin{split}\dot{\bm{\mathrm{I}}}(s,t)=\bm{\mathrm{I}}_{r}(s,t)i+\bm{\mathrm{I}}_{g}(s,t)j+\bm{\mathrm{I}}_{b}(s,t)k\end{split}$
@@
Quaternion Structural Similarity (QSSIM) [28] measures the similarity between two color images in the quaternion domain. It is defined by
$QSSIM({\dot{\bm{\mathrm{X}}}},{\dot{\bm{\mathrm{Y}}}})=\overline{(\dot{a})}\dot{b},$ (1)
@@
$\dot{a}=\frac{2\overline{\dot{\mu}_{\dot{\bm{\mathrm{X}}}}}\dot{\mu}_{\dot{\bm{\mathrm{Y}}}}+C_{1}}{\overline{{\dot{\bm{\mathrm{\mu}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}+C_{1}},$ $\dot{b}=\frac{2\overline{\dot{\sigma}_{\dot{\bm{\mathrm{X}}}}}\dot{\sigma}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}{\overline{{\dot{\bm{\mathrm{\sigma}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}.$ (2)
@@ III Proposed Framework
This section introduces our quaternion multi-focus color image fusion (QMCIF) framework. Section A provides an overview of our QMCIF framework. Section B introduces our quaternion sparse decomposition model and its optimization in detail. Section C presents our quaternion base-detail fusion strategy. Section D presents our quaternion structural similarity refinement strategy. Section E extends our framework to multiple-color image fusion.
@@ III-A Overview
Our QMCIF framework is a patch-wise image fusion framework as shown in Fig. 1. Different from most existing works that perform color image fusion independently, our framework can handle all color channels simultaneously within a single fusion process. We first encode the input color images into their quaternion representations and then use the quaternion sparse decomposition model to jointly learn the optimal coefficient matrix and detail layer for each input image for high-precision focus detection. We employ the quaternion base-detail fusion strategy to fuse the color images in base and detail scales. In the detail-scale fusion process, the two-directional gradients of the detail layers are used to estimate the focus level of each input quaternion representation and select focused patches, generating a detail-scale fused result. In the base-scale fusion process, coefficient matrices are combined with the gradients of the corresponding detail layers to assess the focus level of each input quaternion representation and select focused patches to form the base-scale fused result. To further improve the fusion quality, we design a quaternion structural similarity refinement strategy to merge these fused results patch-by-patch to produce the final fused result using a quaternion weighted structural similarity index measure.
Figure 2: Our quaternion sparse decomposition (QSD) model.
III-B Quaternion sparse decomposition
III-B1 Model formulation
@@
Given a partially focused color image $\bm{\mathrm{I}}$ $\in\mathbbm{R}^{M\times N\times 3}$ and its corresponding quaternion representation $\dot{\bm{\mathrm{I}}}$. Our goal is to decompose texture information from the static background as a strong focus measure. Our quaternion sparse decomposition (QSD) can be formulated as follows.
$\mathop{\arg\min}_{\dot{\bm{\mathrm{D}}},\dot{\bm{\mathrm{Z}}}}\ \ \sum_{k=1}^{K}{\|\dot{\bm{\mathrm{Z}}}_{k}\|}_{*}+\alpha\sum_{n=1}^{2}\|\nabla_{n}\dot{\bm{\mathrm{B}}}\|_{1}$ (3)
$+\beta\|\dot{\bm{\mathrm{D}}}\|_{1}+\lambda\|\dot{\bm{\mathrm{E}}}\|_{F}^{2}$
$s.t.\ \ \dot{\bm{\mathrm{I}}}=\dot{\bm{\mathrm{B}}}+\dot{\bm{\mathrm{D}}}+\dot{\bm{\mathrm{E}}},\ \mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}=\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}_{k}.$
@@
In Eq. (3), ${\nabla}_{1}$ and ${\nabla}_{2}$ represent the vertical and horizontal gradients in the quaternion domain, respectively. The operator $\mathcal{R}(\cdot)$ extracts image patches of size $\sqrt{d}\times\sqrt{d}$ and stacks their vectorized forms into a quaternion matrix in $\mathbbm{H}^{d\times P}$. Column vectors can be divided into $K$ groups. A set of groups is denoted as $\{\mathcal{R}(\dot{\bm{\mathrm{B}}})_{1},\mathcal{R}(\dot{\bm{\mathrm{B}}})_{2},\cdots,\mathcal{R}(\dot{\bm{\mathrm{B}}})_{K}\}$, namely $\mathcal{R}(\dot{\bm{\mathrm{B}}})$ where $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$ represents the $k^{th}$ group. $\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}$ captures the low-rank property of $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$. The nuclear norm of $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ serves as a convex relaxation of the rank function. The positive parameters $\alpha$, $\beta$ and $\lambda$ balance the effects of each term in the optimization process.
@@
As displayed in Fig. 2, our QSD converts a color image to its quaternion representation $\dot{\bm{\mathrm{I}}}$ that is then decomposed into a base-scale layer $\dot{\bm{\mathrm{B}}}$ with low-rank structural information, a detail-scale layer $\dot{\bm{\mathrm{D}}}$ with sparse details and a Gaussian noise layer $\dot{\bm{\mathrm{E}}}$. All layers are quaternion matrices belonging to $\mathbbm{H}^{M\times N}$. To fully utilize the focus information in $\dot{\bm{\mathrm{B}}}$, QSD uses nonlocal similarity to extract local patches in $\dot{\bm{\mathrm{B}}}$, stack these patches into column vectors, and partition them into several groups using the K-means clustering method. QSD learns a set of coefficient matrices $\dot{\bm{\mathrm{Z}}}_{k}$ from these groups using the quaternion dictionary $\dot{\bm{\mathrm{A}}}$ $\in\mathbbm{H}^{d\times L}$ containing $L$ atoms and concatenates these coefficient matrices to obtain $\dot{\bm{\mathrm{Z}}}$ $\in\mathbbm{H}^{L\times P}$. This ensures that patches within the same group are similar.
In general, the detail layer intuitively reflects the sharpness of edge pixels, as it exhibits a greater amplitude of pixel intensity in focused regions. However, in low-texture and boundary areas, the detail layer $\dot{\bm{\mathrm{D}}}$ often provides insufficient features. This may degrade fusion performance. To capture the focus information of smooth regions, coefficient matrix $\dot{\bm{\mathrm{Z}}}$ is introduced as a base-scale measure using quaternion sparse coding [21]. Obtaining the optimal $\dot{\bm{\mathrm{D}}}$ and $\dot{\bm{\mathrm{Z}}}$ simultaneously offers our QSD the greater flexibility than handcrafted filters.
III-B2 Optimization
@@
To solve the optimization problem of our QSD in Eq. (3), we here introduce an iterative learning algorithm under the framework of the quaternion alternating direction method of multipliers [25, 29]. According to the definitions and lemmas in Section II, we write the Lagrangian function of our QSD in Eq. (3) as follows:
@@
[…]
Update ${\{\dot{\bm{\mathrm{J}}}{{}_{k}}\}}_{k=1}^{K}$. Fix ${\{\dot{\bm{\mathrm{Z}}}{{}_{k}}\}}_{k=1}^{K}$ and ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}\}}_{k=1}^{K}$. Let $\dot{\bm{\mathrm{P}}}=$ $\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}}{\mu}$. $\dot{\bm{\mathrm{J}}}{{}_{k}}$ is minimized using the quaternion singular value thresholding method [26].
$\dot{\bm{\mathrm{J}}}{{}_{k}}$ $=\mathop{\arg\min}_{\dot{\bm{\mathrm{J}}}{{}_{k}}}\ \ {\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\frac{\mu}{2}\|\dot{\bm{\mathrm{J}}}{{}_{k}}-(\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}}{\mu})\|_{F}^{2}$ (5)
$=\mathop{\arg\min}_{\dot{\bm{\mathrm{J}}}{{}_{k}}}\ \ \frac{1}{\mu}{\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\frac{1}{2}\|\dot{\bm{\mathrm{J}}}{{}_{k}}-\dot{\bm{\mathrm{P}}}\|_{F}^{2}.$
Therefore, Eq. (5) can be solved using the Lemma 1.
@@
We set the gradient of Eq. (6) with respect to $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ to zero. $\dot{\bm{\mathrm{Z}}}{{}_{k}}$ is explicitly updated as follows.
@@
[…]
The solution to Eq. (8) is obtained using the soft-thresholding method in Lemma 2. Similarly, $\dot{\bm{\mathrm{G}}}{{}_{2}}$ is updated in the same way as the $\dot{\bm{\mathrm{G}}}{{}_{1}}$-subproblem.
@@
$\begin{split}\dot{\bm{\mathrm{B}}}=\mathcal{F}{{}^{-1}}({\frac{\dot{\bm{\mathrm{\Sigma}}}}{\mathcal{F}(\nabla_{1}^{\mathrm{T}}\nabla_{1}+\nabla_{2}^{\mathrm{T}}\nabla_{2})+2}}).\end{split}$ (10)
@@
Eq. (11) can also be solved using Lemma 2.
@@
The optimization process of the QSD model is shown in Algorithm 1.

### v2

# Quaternion Sparse Decomposition for Multi-focus Color Image Fusion

@@ Abstract
Multi-focus color image fusion refers to integrating multiple partially focused color images to create a single all-in-focus color image. However, existing methods struggle with complex real-world scenarios due to limitations in handling color information and intricate focus information. To address these challenges, this paper proposes a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain. This framework introduces 1) a quaternion consistency-aware focus detection method to jointly learn fine-scale details and structure information of color images and generate patch-wise dual-scale focus maps for high-precision focus detection, 2) a quaternion base-detail fusion strategy to obtain dual-scale initial fusion results across input color images, and 3) a quaternion structural similarity refinement strategy to adaptively select optimal patches from initial fusion results and produce the final fused result that preserves fine details and spatial consistency. Extensive experiments demonstrate that the proposed framework outperforms state-of-the-art methods.
@@
Multi-focus color image fusion, quaternion color image fusion, quaternion image decomposition.
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
@@ I Introduction
A single color image typically fails to maintain all scene objects simultaneously in focus due to inherent optical limitations of camera lenses [1]. Specifically, when capturing scenes containing objects at varying depths, only regions within the camera’s focal plane appear sharp while areas outside this plane inevitably suffer from defocus blur. To overcome this limitation, multi-focus color image fusion (MCIF) techniques integrate multiple partially-focused color images of the same scene by employing focus detection and appropriate fusion strategies to produce a single all-in-focus color image [1, 2]. This technology has numerous practical applications, such as digital photography enhancement, preprocessing for subsequent image analysis tasks [3], image segmentation [2] and object recognition [4].
Recent advances in deep learning have significantly promoted MCIF development to enhance multi-scale feature extraction capabilities and improve focused-region discrimination through decision map-based networks [5, 6] and end-to-end architectures [7, 8, 9]. However, it is impractical to acquire ideal ground truth (all-in-focus color images) in real-world conditions [10]. These deep-learning-based methods typically rely on synthetic training datasets for training [11]. Consequently, their generalization capability on real-world images remains limited.
@@
Conventional MCIF methods depend only on human prior knowledge to identify focused objects without the training phase [12]. These methods commonly extend grayscale fusion strategies directly to color images either using grayscale conversion [13] or channel-wise processing [14]. Grayscale-conversion-based MCIF methods first convert the input color images into grayscale images and perform intensity-driven focus detection to obtain focus maps from the grayscale inputs for guiding the fusion process of original color images [13]. For instance, [15] averaged color channels into a grayscale image and obtained focus maps by multi-scale morphological filtering. [16] transformed the color image into the YUV color space and detected focus levels employing the zero crossing and canny edge detectors. However, these methods may overlook multi-channel sharpness variations and cross-channel correlations. This leads to inaccurate focus detection, ghosting artifacts, and blurred boundaries in fused color results [17, 18].
Channel-wise processing-based MCIF methods treat RGB channels independently and apply grayscale fusion methods separately to each channel. For instance, Yang et al. [14] applied sparse coding techniques to MCIF and separately measured focus levels in each channel and fused sparse coefficients channel-by-channel. They fail to consider the cross-channel correlations and introduce spectral-spatial inconsistencies and undesirable color distortions in the fused results [19].
Quaternion representation as a promising tool of color image processing can overcome the aforementioned limitations of color image processing in conventional MCIF methods [20, 21]. It effectively captures color information and preserves cross-channel correlations by treating RGB channels as imaginary components in the quaternion domain [22, 23]. Recently, a quaternion higher-order singular value decomposition (QHOSVD) model [21] was developed by modeling multi-focus color images as third-order quaternion tensors and successfully maintained cross-channel consistency in the stages of focus detection and fusion. However, QHOSVD selects fusion patches based on singular-value energy. This may result in spatial inconsistency and degrade the visual quality of the fused color images.
@@
Motivated by these critical issues, we propose a quaternion multi-focus color image fusion (QMCIF) framework to explicitly designed to simultaneously ensure high-quality multi-focus color image fusion in both spatial and color channels dimensions. Our main contributions are presented as follows:
@@
  • We propose a QMCIF framework to perform color image fusion completely in the quaternion domain. QMCIF is able to obtain high-quality fused color images under various complex scenarios.
  • To effectively measure the focus level and enhance the capability of focus detection in uncertain regions, QMCIF introduces a quaternion consistency-aware focus detection method to explicitly learn focus features of color images and derive dual-scale focus maps in both high-texture and low-gradient regions.
@@
  • To preserve color structure and detail information, QMCIF introduces a patch-wise quaternion base-detail fusion strategy to fuse partially focused color images in detail-scale and base-scale individually.
@@
  • To balance the trade-off between focus information preservation and artifact removal, QMCIF further introduces a quaternion structural similarity refinement strategy to produce the final high-quality fused color image.
@@
  • Extensive experiments on various challenging scenarios demonstrate that QMCIF outperforms the state-of-the-art methods.
@@
The rest of this paper is organized as follows: Section II presents the preliminaries of quaternion algebra. Section III introduces our framework. Section IV presents the experiments and comparisons. Finally, Section V gives the conclusions.
@@ II Preliminaries
[…]
@@
    FrobeniusFrobenius norm: ${\|\dot{\bm{\mathrm{A}}}\|}_{F}={(\sum_{s=1}^{M}\sum_{t=1}^{N}{{|\dot{a}_{s,t}|}^{2}})}^{\frac{1}{2}}$
@@
Given a color image $\bm{\mathrm{I}}$ in three-dimensional real space, its quaternion representation is defined as:
$\begin{split}\dot{\bm{\mathrm{I}}}(s,t)=\bm{\mathrm{I}}_{r}(s,t)i+\bm{\mathrm{I}}_{g}(s,t)j+\bm{\mathrm{I}}_{b}(s,t)k\end{split}$ (1)
@@
Quaternion Structural Similarity ($Q_{SSIM}$) [27] measures the similarity between two color images in the quaternion domain. It is defined by
$Q_{SSIM}({\dot{\bm{\mathrm{X}}}},{\dot{\bm{\mathrm{Y}}}})=\overline{(\dot{a})}\dot{b},$ (2)
@@
$\dot{a}=\frac{2\overline{\dot{\mu}_{\dot{\bm{\mathrm{X}}}}}\dot{\mu}_{\dot{\bm{\mathrm{Y}}}}+C_{1}}{\overline{{\dot{\bm{\mathrm{\mu}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\mu}}_{\dot{\bm{\mathrm{Y}}}}+C_{1}},$ $\dot{b}=\frac{2\overline{\dot{\sigma}_{\dot{\bm{\mathrm{X}}}}}\dot{\sigma}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}{\overline{{\dot{\bm{\mathrm{\sigma}}}}_{\dot{\bm{\mathrm{X}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{X}}}}+\overline{{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}}{\dot{\sigma}}_{\dot{\bm{\mathrm{Y}}}}+C_{2}}.$
@@ III Proposed Framework
This section presents our quaternion multi-focus color image fusion (QMCIF) framework in detail. Section A provides an overview of our QMCIF framework. Section B introduces our quaternion consistency-aware focus detection method. Section C presents our quaternion base-detail fusion strategy. Section D presents our quaternion structural similarity refinement strategy.
@@ III-A Overview
Fig. 1 illustrates our quaternion multi-focus color image fusion (QMCIF) framework. Unlike most existing methods that process color images in a grayscale manner, QMCIF performs joint fusion across all color channels within a unified quaternion representation. Input color images $\bm{\mathrm{I}}_{1}$, $\bm{\mathrm{I}}_{2}$ are converted into their quaternion representations $\dot{\bm{\mathrm{I}}}_{1}$, $\dot{\bm{\mathrm{I}}}_{2}$ using Eq. (1). A quaternion consistency-aware focus detection method jointly estimates their optimal coefficient matrices and detail layers to construct dual-scale focus maps for accurate focus estimation. Under the guidance of dual-scale focus maps, a patch-wise quaternion fusion strategy is applied to fuse the quaternion representations $\dot{\bm{\mathrm{I}}}_{1}$ and $\dot{\bm{\mathrm{I}}}_{2}$ at the base and detail scales to generate the initial fusion results $\dot{\bm{\mathrm{F}}}_{1}$ and $\dot{\bm{\mathrm{F}}}_{2}$. They are further applied with a quaternion structure-similarity-based refinement strategy patch-by-patch to generate the final fused result $\dot{\bm{\mathrm{F}}}$ that is converted back into the real domain to obtain a high-quality all-in-focus color image. Notably, in additional to fuse two color images at a time, our QMCIF framework can be applied to fuse multiple input color images simultaneously. This advantage is verified by the experiments in Figs. 7 and 8.
III-B Quaternion consistency-aware focus detection
To achieve high-precision focus estimation in both high-texture and low-gradient regions, our quaternion consistency-aware focus detection method contains two main steps: (1) a quaternion focal element decomposition model to effectively extract focus-related features in the quaternion domain; and (2) a patch-wise dual-scale focus map generation strategy. Its flowchart is shown in Fig. 2.
Figure 2: Our quaternion consistency-aware focus detection (QCAFD).
III-B1 Quaternion focal element decomposition
@@
Given an input quaternion representation $\dot{\bm{\mathrm{I}}}$ $\in\mathbbm{H}^{M\times N}$, our quaternion focal element decomposition (QFED) model decomposes texture information from its static background as a strong focus measure. Our QFED can be formulated as follows.
$\mathop{\arg\min}_{\dot{\bm{\mathrm{D}}},\dot{\bm{\mathrm{Z}}}}$ $\sum_{k=1}^{K}{\|\dot{\bm{\mathrm{Z}}}_{k}\|}_{*}+\alpha({\|\nabla_{1}\dot{\bm{\mathrm{B}}}\|}_{1}+{\|\nabla_{2}\dot{\bm{\mathrm{B}}}\|}_{1})$ (3)
$+\beta\|\dot{\bm{\mathrm{D}}}\|_{1}+\lambda\|\dot{\bm{\mathrm{E}}}\|_{F}^{2},$
$s.t.\ \ \dot{\bm{\mathrm{I}}}=\dot{\bm{\mathrm{B}}}+\dot{\bm{\mathrm{D}}}+\dot{\bm{\mathrm{E}}},\ \mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}=\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}_{k}$
@@
where ${\nabla}_{1}$ and ${\nabla}_{2}$ represent the quaternion vertical and horizontal gradients respectively. Operator $\mathcal{R}(\cdot)$ extracts image patches of size $\sqrt{d}\times\sqrt{d}$ and stacks their vectorized forms into a quaternion matrix in $\mathbbm{H}^{d\times P}$. Column vectors are divided into $K$ groups denoted as $\mathcal{R}(\dot{\bm{\mathrm{B}}})=\{\mathcal{R}(\dot{\bm{\mathrm{B}}})_{1},\mathcal{R}(\dot{\bm{\mathrm{B}}})_{2},\cdots,\mathcal{R}(\dot{\bm{\mathrm{B}}})_{K}\}$ where $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$ represents the $k^{th}$ group. $\dot{\bm{\mathrm{A}}}\dot{\bm{\mathrm{Z}}}{{}_{k}}$ captures the low-rank property of $\mathcal{R}(\dot{\bm{\mathrm{B}}})_{k}$. $\|\dot{\bm{\mathrm{Z}}}_{k}\|$$_{*}$ serves as a convex relaxation of the rank function. The positive parameters $\alpha$, $\beta$ and $\lambda$ balance the effects of each term in the optimization process.
@@
Our QFED decomposes $\dot{\bm{\mathrm{I}}}$ into a base-scale layer $\dot{\bm{\mathrm{B}}}$ with low-rank structural information, a detail-scale layer $\dot{\bm{\mathrm{D}}}$ with sparse details and a Gaussian noise layer $\dot{\bm{\mathrm{E}}}$. To fully utilize the focus information in $\dot{\bm{\mathrm{B}}}$, QFED uses nonlocal similarity to extract local patches in $\dot{\bm{\mathrm{B}}}$, stacks these patches into column vectors, and partitions them into several groups using K-means clustering method. QFED learns a set of coefficient matrices $\dot{\bm{\mathrm{Z}}}_{k}$ from these groups using pretrained quaternion dictionary $\dot{\bm{\mathrm{A}}}$ $\in\mathbbm{H}^{d\times L}$ and concatenates these coefficient matrices to obtain $\dot{\bm{\mathrm{Z}}}$ $\in\mathbbm{H}^{L\times P}$. This ensures that patches within the same group are similar.
@@
Optimization. To solve the optimization problem of our QFED in Eq. (3), we here introduce an iterative learning algorithm under the framework of the quaternion alternating direction method of multipliers [24]. According to the definitions and lemmas in Section II, we write the Lagrangian function of our QFED in Eq. (3) as follows:
@@
where ${\{{\dot{\bm{\mathrm{J}}}}_{k}\}}_{k=1}^{K}\in\mathbbm{H}^{L\times P}$ represent a set of auxiliary variables to replace ${\{{\dot{\bm{\mathrm{Z}}}}_{k}\}}_{k=1}^{K}$. Matrices ${\dot{\bm{\mathrm{G}}}}_{1}$ and ${\dot{\bm{\mathrm{G}}}}_{2}\in\mathbbm{H}^{M\times N}$ correspond to quaternion horizontal and vertical gradient matrices, replacing $\nabla_{1}\dot{\bm{\mathrm{B}}}$ and $\nabla_{2}\dot{\bm{\mathrm{B}}}$ respectively. Variables ${\{{\dot{\bm{\mathrm{Y}}}}_{1,k}\}}_{k=1}^{K}$ $\in\mathbbm{H}^{L\times P}$, ${\{{\dot{\bm{\mathrm{Y}}}}_{2,k}\}}_{k=1}^{K}$ $\in\mathbbm{H}^{d\times P}$, ${\dot{\bm{\mathrm{Y}}}}_{3},{\dot{\bm{\mathrm{Y}}}}_{4}$ and ${\dot{\bm{\mathrm{Y}}}}_{5}\in\mathbbm{H}^{M\times N}$ are the quaternion Lagrangian multipliers in the optimization process. $\left\langle\cdot\right\rangle$ denotes a quaternion trace product while $\mu$ is a penalty factor. Variables $\dot{\bm{\mathrm{J}}}{{}_{k}}$, $\dot{\bm{\mathrm{Z}}}{{}_{k}}$, $\dot{\bm{\mathrm{Y}}}{{}_{1,k}}$ and $\dot{\bm{\mathrm{Y}}}{{}_{2,k}}$, $\dot{\bm{\mathrm{G}}}{{}_{1}}$, $\dot{\bm{\mathrm{G}}}{{}_{2}}$, $\dot{\bm{\mathrm{D}}}$, $\dot{\bm{\mathrm{E}}}$, $\dot{\bm{\mathrm{Y}}}{{}_{3}}$, $\dot{\bm{\mathrm{Y}}}{{}_{4}}$, $\dot{\bm{\mathrm{Y}}}{{}_{5}}$ are updated in each iteration.
Update ${\{\dot{\bm{\mathrm{J}}}{{}_{k}}\}}_{k=1}^{K}$. Fix ${\{\dot{\bm{\mathrm{Z}}}{{}_{k}}\}}_{k=1}^{K}$ and ${\{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}\}}_{k=1}^{K}$. Let $\dot{\bm{\mathrm{P}}}=$ $\dot{\bm{\mathrm{Z}}}{{}_{k}}+\frac{\dot{\bm{\mathrm{Y}}}{{}_{1,k}}}{\mu}$. The subproblem of $\dot{\bm{\mathrm{J}}}{{}_{k}}$ is reduced to:
$\dot{\bm{\mathrm{J}}}{{}_{k}}=\mathop{\arg\min}_{\dot{\bm{\mathrm{J}}}{{}_{k}}}\ \ \frac{1}{\mu}{\|\dot{\bm{\mathrm{J}}}_{k}\|}_{*}+\frac{1}{2}\|\dot{\bm{\mathrm{J}}}{{}_{k}}-\dot{\bm{\mathrm{P}}}\|_{F}^{2}.$ (5)
Eq. (5) is solved using the Lemma 1 in Section II.
@@
$\dot{\bm{\mathrm{Z}}}{{}_{k}}$ is explicitly updated as follows.
@@
$\begin{aligned} \dot{\bm{\mathrm{G}}}{{}_{1}}=\mathop{\arg\min}_{\dot{\bm{\mathrm{G}}}{{}_{1}}}\ \ \frac{\alpha}{\mu}{\|{\dot{\bm{\mathrm{G}}}}_{1}\|}_{1}+\frac{1}{2}\|\dot{\bm{\mathrm{G}}}{{}_{1}}-\dot{\bm{\mathrm{M}}}_{1}\|{{}_{F}^{2}}\end{aligned}$ (8)
The solution to Eq. (8) is obtained using the soft-thresholding method in Lemma 2 in Section II. Similarly, $\dot{\bm{\mathrm{G}}}{{}_{2}}$ is updated in the same way as the $\dot{\bm{\mathrm{G}}}{{}_{1}}$-subproblem.
@@
[…]
@@
Eq. (11) can also be solved using Lemma 2 in Section II.
@@
The optimization process of the QFED model is shown in Algorithm 1.